OpenStudy (anonymous):

Find the solution set. Separate the two values with a comma. (Problem is in comments)

OpenStudy (anonymous):

\[\LARGE b^2-64=0\]

OpenStudy (mathmale):

First, Dakota, what do you mean by solution set? Second: would this equation be any easier to solve if you were to (1) factor the left side, or (2) move that 64 to the other side by additng 64 to both sides?

OpenStudy (anonymous):

Soltuion set such as .. Hard to explain as I struggle with this subject, but the last answer to an equation was {0,-7}. I would say whichever one you're comfortable with or whichever one is easier.

OpenStudy (mathmale):

Dak: I'd suggest you choose which method would be easier for YOU. That (0,-7) is a point at which x=0 and y=-7. Were you to substitute those values back into the original equation that was given you, the equation would be true, since (0,-7) is a solution.

OpenStudy (mathmale):

So please take a look at the current problem and choose whichever method seems easier for you. Apply it. Try solving for b.

OpenStudy (anonymous):

8 * 8 = 64 however that does not give me the second solution, this is where I'm confused.

OpenStudy (mathmale):

b^2-64=0 could be factored into (b-8)(b+8)=0. This would give you two equations: b-8=0 and b+8=0. What are these two b values?

OpenStudy (mathmale):

Another approach would be to move that 64 to the other side of the equation. Then b^2=64. Taking the square root of both sides, we get b= (plus OR minus) 8. Again, two solutions. Lastly, you're asked to write the "solution set." Remember set notation? If the two solutions are b=8 and b=-8, then the set of solutions is {8,-8}.

OpenStudy (anonymous):

Oh my god, you're a life saver. Thanks so much @mathmale!!! :D

OpenStudy (mathmale):

My great pleasure! :)